40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4846     Entry time: Tue Jun 21 00:38:21 2011
Author: Sonali 
Type: Update 
Category: Green Locking 
Subject: repositioned "QPDY_PD" 

1.The aim is the laser frequency stabilisation of PSL and AUX.

2.As a first step we want to couple some of the AUX laser beam into a single mode optical fibre and route the fibre to the PSL table.

3.The position of the optical fibre on the ETMY table is shown by the coupler in the attached picture. The yellow lines show the new scheme we want to implement.

4.WHAT WE DID TODAY.

  • The Y-arm was locked so that we could use the transmitted IR beam as the reference.
  • We shifted the position of the "QPDY_PD" .
  • We also shifted the "ETMYT" camera to make space for the "QPDY_PD".
  • The mirror  directing the beam into the "QPDY_PD" was rotated by 90 degrees to adhere to the new position of the "QPDY_PD".
  •  The attached photo shows the table as it is right now after the repositioning.
5.We continue with the positioning of the fibre-coupling tomorrow.
Attachment 1: ETMY_table_2011_June20.png  3.271 MB  | Hide | Hide all
ETMY_table_2011_June20.png
Attachment 2: ETMY_table_change1.jpg  3.249 MB  | Hide | Hide all
ETMY_table_change1.jpg
ELOG V3.1.3-